OpenStudy (anonymous):

John is twice as old as Mary. The sum of their ages is 21. How old is Mary? Let J = John's age and M = Mary's age. Select the system equations that represents the problem. a)J - 2M = 0 b)M - 2J = 0 c)J + M = 21 d)J = M + 2 There can be more than one answer :)

OpenStudy (shadowfiend):

So first off, is there an equation that expresses that the sum of the ages is 21?

OpenStudy (bahrom7893):

J = 2M=> =>J - 2M = 0 J + 2M = 21 (sum of both is 21) So: J - 2M = 0 J+2m = 21